OpenStudy (anonymous):

A patio in the shape of a semicircle has a diameter of 21 ft. What is the approximate area of the patio, rounded to the nearest square foot? Use 3.14 to approximate pi .

OpenStudy (anonymous):

@Xnlovely

OpenStudy (anonymous):

10.5*10.5=110.25 110.25*3.14=346.185 346.185/2=173

OpenStudy (anonymous):

you copied that from someone elses didnt chu :3

OpenStudy (anonymous):

no i actually worked that one out

OpenStudy (anonymous):

Lol dun worry

OpenStudy (anonymous):

i swear i didnt even go away from this site lol

OpenStudy (anonymous):

but whatever :P